A THEORETICAL STUDY OF FeNC IN THE ^{6}Δ ELECTRONIC GROUND STATE

\underline{TSUNEO HIRANO}, REI OKUDA, and UMPEI NAGASHIMA, Research Institute for Computational Sciences, National Institute of Advanced Industrial Science and Technology, 1-1-1 Umezono, Tsukuba, Ibaraki 305-8568, Japan; VLADIMÍR SPIRKO, Center for Biomolecules and Complex Molecular Systems, Institute of Organic Chemistry and Biochemistry, Academy of Sciences of the Czech Republic, Flemingovo nám. 2, CZ-166 10 Praha 6, Czech Republic; PER JENSEN, Theoretische Chemie, Bergische Universität, D-42097 Wuppertal, Germany.

We report an ab initio calculation, at the MR-SDCI+Q+E_{rel}/[Roos ANO (Fe), aug-cc-pVQZ (C, N)] level of theory, of the potential energy surface for ^{6}Δ_{i} FeNC. From the ab initio results, we have computed values for the standard spectroscopic parameters of FeN^{12}C and FeN^{13}C. Analytical representations of the potential energy surfaces have been fitted through the ab initio points, and the resulting functions have been used for directly solving the rotation-vibration Schrödinger equation by means of the MORBID program and by means of an adiabatic-separation method. For ^{6}Δ_{i} FeNC, our ab initio calculations show that the equilibrium structure is linear with r_{e}(Fe-N) = 1.9354 Å and r_{e}(N-C) = 1.1823 Å. We find that the bending potential is very shallow, and the MORBID calculations show that the zero-point averaged structure is bent with the expectation values ⟨r(Fe-N)⟩ = 1.9672 Å, ⟨r(N-C)⟩ = 1.1866 Å, and ⟨ρ̄⟩ = 180^{∘} − ⟨^{}(Fe-N-C)⟩ = 13^{∘}. The experimentally derived bond length r_{0}(N-C) = 1.03(8) Å reported for ^{6}Δ_{i} FeNC by J. Lie and P. J. Dagdigian [J. Chem. Phys. 114, 2137-2143 (2001)] is much shorter than the corresponding ab initio r_{e}-value and the averaged value from MORBID. Our calculations suggest that this discrepancy is caused by the inadequate treatment of the large-amplitude bending motion of ^{6}Δ_{i} FeNC. It would appear that for floppy triatomic molecules such as FeNC, r_{0}-values have little physical meaning, at least when they are determined with the effects of the large-amplitude motion being ignored, i.e., under the assumption that the r_{0} structure is linear.